### v1

# Multitask and Multimodal Neural Tuning for Large Models

Abstract
In recent years, large-scale multimodal models have demonstrated impressive capabilities across various domains. However, enabling these models to effectively perform multiple multimodal tasks simultaneously remains a significant challenge. To address this, we introduce a novel tuning method called neural tuning, designed to handle diverse multimodal tasks concurrently, including reasoning segmentation, referring segmentation, image captioning, and text-to-image generation. Neural tuning emulates sparse distributed representation in human brain, where only specific subsets of neurons are activated for each task. Additionally, we present a new benchmark, MMUD, where each sample is annotated with multiple task labels. By applying neural tuning to pretrained large models on the MMUD benchmark, we achieve simultaneous task handling in a streamlined and efficient manner. All models, code, and datasets will be publicly available after publication, facilitating further research and development in this field.
Multitask and Multimodal Neural Tuning for Large Models
Hao Sun1, Yu Song2, Jihong Hu2, Yen-Wei Chen2, Lanfen Lin1, 1College of Computer Science and Technology, Zhejiang University 2College of Information Science and Engineering, Ritsumeikan University Email1: {sunhaoxx, llf}@zju.edu.cn Email2: {yusong@fc, gr0609ik@ed, chen@is}.ritsumei.ac.jp
1 Introduction
With the significant progress of large language models (LLMs), researchers have proposed combining large models from multiple modalities to handle multimodal tasks, such as referring segmentation and text-to-image synthesis Ren et al. (2023); Ho et al. (2020). However, these methods primarily focus on individual tasks by designing sophisticated structures or strategies. While this approach can enhance performance in specific areas, the specialized architectural design may limit their applicability to other tasks and domains. Although it is possible to expand the capabilities of LLMs to handle more tasks by developing more intricate structures or tuning strategies, this approach faces scalability challenges.
Figure 1: The overview of our proposed multitask neural tuning. All inputs and outputs are token-based, encompassing both texts and images. The model generates specific tokens for different tasks, such as <OBJ> for segmentation and <GEN> for text-to-image generation. During tuning, a new sparse task network is introduced to emulate SDR and provide task guidance for pretrained LLMs. The entire LLM remains frozen, with only the newly introduced parameters being tunable.
Humans inherently possess multitask learning abilities and can easily adapt to new tasks. One reason for this is the principle known as Sparse Distributed Representation (SDR) or the Sparse Coding Hypothesis (SCH), which describes a way of representing data where only a small fraction of neurons are active at any given time Ahmad and Hawkins (2016); Spanne and Jörntell (2015). This principle is thought to contribute to efficient information processing and the capacity of unique representations in the brain. Inspired by these findings in brain science, we propose a new tuning strategy for multimodal multitask learning, termed neural tuning (as illustrated in Figure 1). Our neural tuning strategy introduces two novel concepts. Firstly, we formulate different tasks in an all-in-token manner, allowing new tasks to be introduced by adding new task tokens, thereby ensuring scalability for additional tasks or modalities. Secondly, we introduce a new sparse task network to emulate SDR for multitask tuning, where only a subset of neurons is activated for each task. The neurons activated by different tasks include both intersecting and unique parts: the intersecting parts represent common information required for various tasks, while the unique parts represent the distinct characteristics of each task. In the design of neural tuning, we avoid additional cross-attention mechanisms and complex decoders to ensure simplicity and efficiency.
However, there are few datasets designed for multimodal multitask learning scenarios, particularly those involving reasoning segmentation tasks where multiple objects can only be segmented through complex reasoning based on both images and text. To address this gap and advance research in this area, we propose a new dataset called MMUD. Constructed using GPT-4 and further annotated by humans, MMUD contains over 36,000 samples. Each sample includes an image with an annotated content caption, a complex reasoning question-answer pair, and referring segmentation masks. We fine-tune pretrained LLMs with neural tuning on MMUD for four tasks: vanilla referring segmentation, reasoning segmentation, image captioning, and text-to-image generation. Figure 1 the demonstrates the pipeline of our proposed multitask neural tuning and the results show that we can achieve state-of-the-art performance.
In summary, our contributions are threefold:
  • We propose neural tuning, a novel framework that unifies diverse multimodal tasks through a concise all-in-token methodology. This innovative approach simplifies the integration of new tasks by merely introducing new task-specific tokens, thus enhancing the flexibility and scalability of large multimodal models.
  • Our method incorporates a sparse task tuning network designed to emulate SDR for various tasks. This network adaptively activates specific neurons to efficiently manage and differentiate between multiple tasks, improving both the precision and adaptability of task performance.
  • We introduce a comprehensive new multimodal benchmark, MMUD, wherein each sample is meticulously annotated for multiple tasks. By fine-tuning models on MMUD, we demonstrate the capability to achieve superior multitask and multimodal task processing performance.
2 Related Works
In this section, we analysis some recent works related to our works, including multimodal tuning, referring segmentation, and text-to-image synthesis.
(a) Detailed structure of our proposed neural tuning in self-attention.
(b) The visualization of the activated neurons in sparse vector.
Figure 2: The detail of proposed neural tuning and the illustration of activated neurons in sparse vector. Figure (a) illustrates the integration of the sparse task network designed to generate task guidance for pretrained LLMs. At the core of the sparse task network is the sparse vector $Z_{k}$, with only a percentage of neurons activated to perform specific tasks. Sparse vectors across different layers are interconnected through an EMA updating mechanism. Figure (b) visualized the sparse vector $Z_{k}^{a}$ for different tasks ($D_{z}$ is set to 128 in the example).
2.1 Multimodal Tuning for Large Models
Recently, many multimodal tuning strategies have been proposed based on language models, aiming to equip LLMs with multimodal perception capabilities. For instance, Flamingo Alayrac et al. (2022) employs a frozen image encoder and integrates multimodal signals using gated cross-attention, demonstrating the potential of LLMs to understand multiple modalities. Then BLIP-2 Li et al. (2023a) introduces the Q-Former module, a sophisticated structure designed to align image and text embeddings by extracting common semantics between modalities. This approach has gained widespread adoption due to its effectiveness. Additionally, FROMAGe Koh et al. (2023) employs linear transformations to ground text features in the visual domain, facilitating seamless translation between text and images. Nevertheless, recent vision-language LLMs (LVLMs) have shown great capability in multimodal understanding, such as LLaVA Liu et al. (2023) and Qwen-VL Bai et al. (2023). Our proposed neural tuning strategy is also based on multimodal tuning of LLMs; however, it places greater emphasis on multitask learning in a more elegant and concise manner (especially for reasoning segmentation under complex scenarios) .
2.2 Referring Segmentation and Reasoning Segmentation
Referring segmentation aims to segment objects in an image based on instructive texts. In earlier years, researchers typically employed a text encoder along with a U-Net-like Ronneberger et al. (2015) vision backbone to generate segmentation masks, such as LAVT Yang et al. (2022) and SLViT Ouyang et al. (2023). With the rapid development of large models, simple referring segmentation tasks cannot adequately verify the multimodal reasoning capabilities of these models. To address this challenge, complex reasoning segmentation is proposed which involve asking relatively complex questions about images that require reasoning based on the image content to respond. When the model answers the question, it should also segment the objects involved. For example, Lai et al. (2023) first designed a new LISA model and introduced a new token for segmentation. PixelLM Ren et al. (2023) was then proposed with a newly designed codebook for multi-instance segmentation.
2.3 Text-to-Image Synthesis
There are two text-to-image synthesis approaches highly related to our work: vector quantized generative adversarial network (VQGAN) related methods Esser et al. (2021) and diffusion-based methods Ho et al. (2020); Song et al. (2020). VQGAN aims to map images into a discrete latent space while diffusion models simulates a diffusion process, where data is gradually transformed from a simple prior distribution (like Gaussian noise) to the complex target distribution. State-of-the-art text-to-image models such as GLIDE Nichol and Dhariwal (2021), DALL-E Ramesh et al. (2021), and Imagen Saharia et al. (2022) leverage advanced techniques in VQGAN and diffusion models to generate high-quality, diverse images. In our work, we primarily employ VQGAN for synthesizing images, but we also explore the potential of combining our approach with diffusion networks for image generation.
3 Method
The overall pipeline of our proposed neural tuning is illustrated in Figure 1. We introduce a new sparse task network into pretrained models, serving as the SDR or SCH. The entire model takes image and text inputs and produces corresponding outputs for different tasks, such as multi-instance referring segmentation (<OBJ>) and image generation (<GEN>). All input and output data are in token format. During the process, the sparse task network activates only a certain percentage of neurons for different tasks. In this section, we first introduce the detailed design of our proposed neural tuning and then provide some details for multitask training.
3.1 Neural Tuning for Large Models
For a multimodal input comprising an image and a sentence, we first embed the text into $I_{txt}=\{I_{txt}^{l}\}_{l=1}^{L_{t}}$, where $L_{t}$ represents the textual length. Subsequently, we utilize a pretrained vision encoder to extract features from the image, resulting in $I_{img}=\{I_{img}^{l}\}_{l=1}^{L_{i}}$, with $L_{i}$ patches. To integrate the multimodal input, we concatenate the text embeddings and image features (if the task requires visual input) to form inputs for the pretrained LLM: $I=[I_{img};I_{txt}]$.
To fine-tune the pretrained LLMs, we introduce a sparse task network. It is parallel to pretrained LLMs but gets linked in each layer for task guidance. For each tuning layer, it maintains a learnable vector named the Sparse Vector and the details are shown in Figure 2(a). In the $k$-th layer of the LLM before self-attention, we first project the hidden embeddings $H$ into a subspace to obtain the sparse vector:
$Z_{k}=W_{z}H\in R^{(L_{t}+L_{i})\times D_{z}},$ (1)
where $H\in\mathbb{R}^{(L_{t}+L_{i})\times D_{h}}$, $W_{z}\in\mathbb{R}^{D_{z}\times D_{h}}$ is a learnable matrix, $D_{h}$ is the embedding size of the pretrained LLM, and $D_{z}$ is the dimension of the sparse vector. To ensure the sparse vector is transmitted between layers, $Z_{k}$ is updated using an Exponential Moving Average (EMA) mechanism with a hyperparameter $\alpha$:
$Z_{k}^{u}=\alpha Z_{k-1}^{u}+(1-\alpha)Z_{k},$ (2)
where $Z_{k-1}^{u}$ is the sparse vector from the last layer. This approach allows the sparse vector to not only receive information from the previous layer but also be aware of the current layer’s information. To enable sparse distributed representation during model flow, only a subset of the sparse vector’s nodes are activated for various tasks. We first randomly sample an activation rate $r$ from a normal distribution $p(r)=N(r;\beta,(0.1\beta)^{2})$, where $\beta\in(0,1)$ is a predefined hyperparameter [^1] More activation patterns are discussed in Appendix C. Then, we activate corresponding neurons as follows:
$Z_{k,j}^{a}=\left\{\begin{matrix}\begin{aligned} &Z_{k,j}^{u}\quad&Z_{k,j}^{u}>=Z_{k,r}^{u}\\ &0\quad&Z_{k,j}^{u}<Z_{k,r}^{u}\end{aligned}\end{matrix}\right.$ (3)
where $Z_{k,r}^{u}$ is the largest top $r$ values and $j\in[1,D_{z}]$. $Z_{k}^{a}$ represents the activated neurons for a specific task. Next, to allow the pretrained LLM to leverage task-specific guidance, we use a linear transformation to project the activated sparse vector back to the LLM’s hidden space:
$Z_{k}^{b}=W_{b}Z_{k}^{a}\in\mathbb{R}^{(L_{t}+L_{i})\times(2D_{h})},$ (4)
where $W_{b}$ is a learnable parameter. For the self-attention mechanism of the LLM, we split $Z_{k}^{b}$ into two parts and use the residual for query ($Q$) and value ($V$) computation,
$Q^{\prime}$ $=W_{q}H+Z_{k}^{b}[:,0:D_{h}],$ (5)
$V^{\prime}$ $=W_{v}H+Z_{k}^{b}[:,D_{h}:2D_{h}],$
$K$ $=W_{k}H.$
$Q^{\prime}$, $K$, and $V^{\prime}$ are then employed for the vanilla self-attention in pretrained LLMs. Overall, the pretrained LLM is responsible for the main inference, while the sparse task network handles the execution of different tasks. Interestingly, our experiments reveal that more related tasks tend to activate a greater number of the same neurons (e.g., vanilla referring segmentation and reasoning segmentation). This finding highlights the relationship between different tasks, as illustrated in Figure 2(b).
In the LLM output stage, we introduced new tokens for multimodal tasks alongside the original textual tokens. For example, we introduced <OBJ> tokens for segmentation and <GLB> tokens for text-to-image synthesis. Different task-specific tokens are fed into their respective decoders. This approach allows us to unify the input and output of all tasks into an all-in-token manner, making it convenient to add more tasks and modalities.
3.2 Multitask Training
As different tasks are unified into an all-in-one token manner, neural tuning can be conducted using a simple causal language modeling approach. Cross-entropy loss is employed as the loss function ($\mathcal{L}_{txt}$) for next-token prediction.
For segmentation-related tasks, the embeddings corresponding to segmentation tokens <OBJ>, $H_{seg}$, are collected and fed to a lightweight decoder for corresponding mask generation:
$\hat{y}=Decoder_{seg}(H_{seg},W_{seg}),$ (6)
where $W_{seg}$ represents the learnable parameters in the mask decoder. Different from previous methods Lai et al. (2023), which employ pretrained SAM Kirillov et al. (2023) for predicting masks, the designed decoder is much lighter. Since the models can generate multiple segmentation tokens at a time, we can easily perform multi-instance segmentation. Following previous works Lai et al. (2023); Ren et al. (2023), DICE loss Milletari et al. (2016) is employed in our framework to guide the segmentation tasks:
$\mathcal{L}_{seg}=\mathcal{L}_{DICE}=\frac{1}{N}\sum_{i}(1-\frac{2|\hat{y_{i}}\cap y_{gt,i}|}{|\hat{y_{i}}|+|y_{gt,i}|})$ (7)
where $N$ is the number of samples, $\hat{y}$ is the predicted mask and $y_{gt}$ is the ground truth.
In text-to-image synthesis tasks, we employ pretrained VQGAN generators Esser et al. (2021); Wang et al. (2022) to synthesize images. To generate a sequence of indices for VQGAN to produce images, we train a conditional transformer to predict the indices in an autoregressive manner:
$p(\mathbf{s}|c)=\prod_{i}p(s_{i}|s_{<i},c),$ (8)
where $\mathbf{s}$ is the sequence of indices for VQGAN to generate images and $c$ is the condition that controls the contents of the images. To simplify the tuning process, we first pretrain the conditional transformer on MS-COCO Lin et al. (2014) and employ the image embeddings from the pretrained CLIP model Radford et al. (2021) as the condition. After training, we can generate images via the CLIP embeddings. Therefore, in neural tuning, we simply align the embeddings of the synthesis tokens ($H_{gen}$) to the CLIP embeddings ($H_{clip}$). The mean squared error is employed to perform the alignment:
$\mathcal{L}_{gen}=\mathcal{L}_{MSE}=\frac{1}{N}\sum_{i}(W_{gen}H_{gen}-H_{clip})^{2},$ (9)
where $W_{gen}$ is a learnable transformation to project token embeddings into the CLIP feature space.
The pretrained LLMs are then tuned with all tasks involved. The overall training loss is represented as:
$\mathcal{L}=\mathcal{L}_{txt}+\lambda_{seg}\mathcal{L}_{seg}+\lambda_{gen}\mathcal{L}_{gen},$ (10)
where $\lambda_{seg}$ and $\lambda_{gen}$ are coefficients to balance the numerical scales of different losses.
Figure 3: Pipeline for MMUD Dataset Generation. First, GPT-4v is used to generate captions describing the image contents and reasoning question-answer pairs. Then to ensure dataset quality, meaningless cases are filtered out. Finally, <OBJ> tokens are manually appended to objects in the answers to help large models better understand the relationships between images and text.
4 MMUD Benchmark
To enable LLMs with multitask and multimodal processing capabilities, a high-quality dataset is essential. However, there are few datasets that provide multitask annotations specifically designed for large models. To address this, we have constructed the MMUD dataset. MMUD dataset contains over 36,000 samples. We have divided the dataset into training, validation, and test subsets, comprising 33,682, 1,400, and 1,400 samples, respectively. Each sample in the dataset is annotated with referring segmentation masks, complex reasoning question-answer pairs, and image captions. In this work, we focus on four tasks with MMUD: vanilla referring segmentation, reasoning segmentation, image captioning and text-to-image generation.
We constructed our dataset using open-source datasets: RefCOCO Yu et al. (2016), RefCOCOg Mao et al. (2016), and RefClef Nagaraja et al. (2016); Mao et al. (2016). Since these datasets only provide referring segmentation masks, we augmented them with annotations for other tasks. The dataset generation pipeline is illustrated in Figure 3. Initially, we employed GPT-4v to generate image captions describing the contents. Subsequently, GPT-4v was used to generate complex questions along with answers based on the image contents. We then performed manual filtration to ensure that the generated contents were meaningful and suitable for multimodal understanding. We filtered out samples in cases where: 1) the generated contents were meaningless; 2) the length of the generated contents was either too long or too short; 3) the answers could be directly inferred from the questions without reference to the images. To enhance the capability of complex reasoning, we manually inserted the <OBJ-i> token after each object in the answer, where $i$ represents the $i$-th object in the image. For example, when generating the question Can you judge the relationship between the two people in the front based on what happened in the picture?" with the answer with answer In this image, one of the women in the front is cutting a pizza, and a little girl is next to her…, we manually inserted the <OBJ> token after the expression of each object: In this image, one of the women in the front <OBJ-1> is cutting a pizza <OBJ-2>, and a little girl <OBJ-3> is next to her…. Following the structure of the RefCOCO datasets, each sample image contains multiple meaningful objects with corresponding masks. Each <OBJ-i> token not only corresponds to an object but also to the corresponding segmentation mask, which could help models better capture the relationship between texts and images. The dataset will be made publicly available to advance research in this area.
5 Experiments and Analysis
5.1 Experimental Settings
In our method, we employ the pretrained LLaMA2-13B and LLaMA2-7B as the textual foundation models. For image feature extraction, we utilize CLIP-ViT-L/14. The images are resized to 224 pixels as inputs. All parameters in LLaMA and CLIP are kept frozen, and only the newly introduced tokens, sparse task network, and task decoders are trainable (2.9% in total). The efficient parameter $\alpha$ in sparse vector updating is set to 0.9, while the neuron activation rate $\beta$ is set to 0.4. We set the dimension of the sparse vector $Z$ to 128. During optimization, to balance the numerical scales in the loss functions, we set $\lambda_{seg}$ to 1.0 and $\lambda_{gen}$ to 10.0. We train the models for 10 epochs with a batch size of 12 and a cosine learning rate decay scheduler. The tuning process takes about 36 hours on four RTX 4090 GPUs or three NVIDIA A100 GPUs.
Table 1: The results our neural tuning (NT in table) and previous methods on MMUD and three public datasets for reasoning segmentation and referring segmentation.
(a) The results of reasoning segmentation on MMUD dataset. As MMUD is a new dataset, we re-implement the methods based on their open-source codes.
Method w/ LLM Valid Test
mIoU oIoU mIoU oIoU
LVAT 21.2 20.2 23.3 23.1
LISA-7B $\checkmark$ 60.1 59.9 61.3 61.8
PixelLM-7B $\checkmark$ 61.1 60.7 63.2 62.6
NT-7B(Ours) $\checkmark$ 62.2 61.6 63.1 62.8
LISA-13B $\checkmark$ 62.0 61.2 63.1 62.7
PixelLM-13B $\checkmark$ 63.4 62.8 64.4 64.0
NT-13B(Ours) $\checkmark$ 63.4 63.0 64.9 64.2
(b) The results of vanilla referring segmentation on RefCOCO, RefCOCO+, and RefCOCOg are presented. The metric used in the table is oIoU.
Method w/ LLM TFLOPs RefCOCO RefCOCO+ RefCOCOg(U)
Val TestA TestB Valid TestA TestB Valid Test
LVAT 72.7 75.8 68.8 62.1 68.4 55.1 61.2 62.1
LISA-7B $\checkmark$ 7.16 74.0 76.3 70.4 62.5 66.3 56.0 67.0 69.1
PixelLM-7B $\checkmark$ 3.57 73.0 76.5 68.2 66.3 71.7 58.3 69.3 70.5
NT-7B(Ours) $\checkmark$ 2.47 74.2 76.7 68.0 66.3 71.2 58.1 70.3 70.2
5.2 Results on MMUD
Table 1 and Table 2 presents the results of our proposed neural tuning compared with previous methods. Some visualized cases are shown in Figure  4 (Appendix A). For vanilla referring segmentation and reasoning segmentation, we mainly compare our method with LVAT, LISA, and PixelLM. We employ the mean intersection over union (mIoU) and overall IoU as the metrics. LVAT aims to fuse the BERT features into vision backbones, achieving great results in vanilla segmentation. However, when it comes to complex reasoning scenarios, LVAT fails to converge while other LLM-based methods yield good results, demonstrating the powerful reasoning capabilities of large models. Compared with LISA and PixelLM, which employ pretrained LLaVA Liu et al. (2023) and focus on complex reasoning segmentation, neural tuning can achieve better results, demonstrating its effectiveness. Regarding inference speed, our tuned 7B model can complete the reasoning segmentation process within 80ms per text-image pair on an RTX 4090 GPU, while the 13B model takes 110ms on an RTX A6000 GPU. This demonstrates the efficiency of our proposed neural tuning. We believe this efficiency results from the linear sparse vector updating and task guidance. Furthermore, to further evaluate the effectiveness of our proposed method, we also present the performance on the original test set of RefCOCO, RefCOCO+, and RefCOCOg datasets  [^2] For fair comparision, we re-split the train, validation, and test set of MMUD according to the original datasets.. As shown in Table 1(b), our method achieves state-of-the-art performance on the public datasets but with lower computational burden, revealing the effectiveness and efficiency of our proposed multitask neural tuning for segmentation tasks.
Table 2: The results our neural tuning and previous methods on MMUD for image captioning task and text-to-image generation task.
Image Captioning Text-to-image Synthesis
Method BLEU-4 METEOR CIDEr Method FID($\downarrow$) KID($\downarrow$) IS
BLIP2-6.7B 41.9 34.7 133.0 GLIGEN 12.9 12.5 31.1
ExpansionNetV2 41.1 34.0 132.7 U-ViT-S/2 13.7 15.9 29.8
mPLUG 43.0 34.1 134.0 Parti 11.0 13.6 30.6
NT-7B(Ours) 43.7 35.5 133.2 NT-7B(Ours) 12.7 15.2 31.4
Table 3: The ablation study on MMUD test set for reasoning segmentation. The LLaMA2-7B is employed in the ablation. $Z_{k}^{U}$ Upd. means the sparse vector updating module (Equation 2) and SDR indicates the sparse distributed representation emulation. Task V.S., Task I.C. and Task I.S. means the vanilla referring segmentation, image captioning and text-to-image synthesis tasks, respectively.
Ablations Tuning Modules MMUD
$Z_{k}^{U}$ Upd. SDR Task V.S. Task I.C. Task I.S. Q K V mIoU oIoU
$\checkmark$ $\checkmark$ $\checkmark$ 57.7 57.6
$\checkmark$ $\checkmark$ $\checkmark$ 59.3 58.2
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 60.0 58.9
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 61.1 50.7
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 62.0 61.2
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 62.9 62.6
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 63.1 62.8
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 62.0 61.1
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 63.5 63.0
For image captioning, we compare our method with BLIP-2 Li et al. (2023a), ExpansionNetV2 Hu et al. (2022) and mPLUG Li et al. (2022). BLIP-2 proposed a Q-Former for multimodal interaction whlie mPLUG learn the relationship between modalities by cross-modality skip-connection. We employ the BLEU-4 Papineni et al. (2002), CIDEr Vedantam et al. (2015), and METEOR Banerjee and Lavie (2005) metrics for evaluation. The results are shown in Table 2. As we can draw from the results, we can reach the competitive performance compared with previous state-of-the-art approaches. On text-to-image synthesis tasks, quantitative metrics are shown in Table 2. We employ FID score Heusel et al. (2017), KID Score Bińkowski et al. (2018), and inception score (IS Salimans et al. (2016)) as the metrics. We compare our method with previous methods, including Parti Yu et al. (2022), GLIGEN Li et al. (2023b), and U-ViT-S/2 Bao et al. (2023). Compared with other methods, one of the significant advantage of our method is that neural tuning is designed for multitask tuning instead of a certain task.
5.3 Ablation Study
To prove that each module of our proposed method is effective, we conducted ablation experiments on complex reasoning segmentation tasks. The ablation of different aspects are as follows:
Modules in neural tuning: The quantitative results are presented in Table 3. We observed that removing sparse vector updating (Equation 2), which renders the tuning layers of the sparse task network independent of each other, leads to a drop in performance. This highlights the significance of interlinking different layers of tuning networks, a factor overlooked by prior tuning methods such as LoRA Hu et al. (2021) and (IA)3 Liu et al. (2022). Additionally, performance degrades when the SDR strategy is eliminated, indicating that all neurons are activated for all tasks. Finally, we conducted an ablation study on multitask learning. The results show that removing the multitask learning strategy led to a relative decrease in performance. Notably, the vanilla referring segmentation and image captioning tasks appear to contribute more to complex reasoning than the image generation tasks, which corresponds to the visualized results in Figure 2(b).
The selection of tuning modules: In Section 3.1, we integrated task guidance into the LLMs for query and value, following established methods Hu et al. (2021). However, this approach is not the sole option. We conducted corresponding ablation studies, with the quantitative results presented in Table 2. The results indicate that integrating task guidance into keys and values tends to result in poorer performance. While performance improves when task guidance is integrated into all query, key, and value modules, the improvement is only marginal compared to when it is integrated solely into query and value. Therefore, to strike a balance between performance and complexity, we opted to perform tuning only on query and value modules (Equation 5).
Choice of image generation decoders: Apart from VQGAN, there are alternative options for image decoding. We also utilized a pre-trained DALL-E2 decoder [^3] http://github.com/hidden/for/review for generating images from the hidden embeddings. A comparison is presented in Figure A. The quality and style of the generated images are significantly influenced by the pre-trained decoders. This bias is a result of the pretraining data used for the decoders. Hence, in theory, neural tuning can employ any pre-trained image synthesis decoder. Unfortunately, due to hardware limitations, we were unable to test other pre-trained models for image generation.
6 Conclusion
In this paper, we introduce a novel tuning strategy called neural tuning. Neural tuning unifies tasks into an all-in-token approach, enhancing scalability for additional modalities or tasks. Furthermore, we emulate human thought processes using sparse distributed representation by activating specific neurons for different tasks. We evaluate our method across four tasks, including reasoning segmentation and text-to-image synthesis, demonstrating competitive performance compared to current state-of-the-art methods. To further research in this field, we introduce the MMUD dataset, which includes various annotations for different tasks.
7 Limitations
One of the biggest limitation is that our research currently does not incorporate acoustic tasks into neural tuning and MMUD. Ideally, under optimal conditions, neural tuning could accommodate all types of multimodal tasks by appropriately formulating them into tokens. In the future, we aim to conduct more studies on multitask learning for large multimodal models, exploring new technologies, models, and benchmarks. Additionally, similar to other large-scale models, this work carries the potential risk of misuse. To mitigate this issue, we will implement an application-based access system for the release of data and models. This approach aims to ensure their use remains within reasonable and ethical boundaries.
Appendix A Case Visualization
Figure 4: The visualization of some cases from our model, including vanilla referring segmentation, reasoning segmentation, image captioning, and image generation.
The visualization of some cases from our model are shown in Figure A. The tasks included in these cases include: vanilla referring segmentation, reasoning segmentation, image captioning, and image generation. For image generation, we employ two pretrained decoders: VQGAN and DALL-E2.
Table 4: The ablation of $\beta$ in sparse vector updating.
Tasks involved oIoU of Reasoning Segmentation on MMUD
$\beta=0.1$ $\beta=0.2$ $\beta=0.3$ $\beta=0.4$ $\beta=0.5$ $\beta=0.6$ $\beta=0.7$ $\beta=0.8$ $\beta=0.9$ $\beta=1.0$
4 Tasks 62.0 61.9 62.5 62.8 62.1 61.6 60.7 61.0 60.6 61.0
3 Tasks 60.4 60.9 60.7 61.1 62.5 62.5 61.8 61.9 61.2 60.8
2 Tasks 60.0 60.3 60.4 61.2 60.6 61.9 61.7 61.9 62.2 62.0
1 Task 60.1 60.4 59.8 60.7 60.3 60.9 61.5 61.3 61.6 61.7
Appendix B Ablation of $\beta$ in Sparse Vector Updating
The detailed ablations regarding $\beta$ in sparse vector updating are presented in Table 4. In these experiments, we consistently use the reasoning segmentation task and employ the corresponding oIoU for evaluation. For instance, two tasks involved refer to reasoning segmentation and vanilla segmentation, while only one task involved refers to reasoning segmentation only. From the results, we observe that setting $\beta$ to 0.4 yields the best performance when all four tasks are involved. For scenarios involving three tasks, the optimal performance is achieved with $\beta$ set to 0.5 or 0.6. Finally, when there are only one or two tasks, it is more beneficial to activate a higher proportion of neurons, with $\beta$ set to 0.9 or 1.0.
Table 5: The ablation of activation patterns for sparse vector. In the experiments, all of four tasks are employed.
Activation Pattern oIoU of Reasoning Segmentation on MMUD
$\beta=0.1$ $\beta=0.2$ $\beta=0.3$ $\beta=0.4$ $\beta=0.5$ $\beta=0.6$ $\beta=0.7$ $\beta=0.8$ $\beta=0.9$ $\beta=1.0$
Gaussian Random 62.0 61.9 62.5 62.8 62.1 61.6 60.7 61.0 60.6 61.0
Top-$2\beta$ Random 61.8 62.0 61.9 62.9 62.7 61.2 61.0 60.9 60.5 59.7
Level Random 61.8 61.7 62.2 62.6 62.6 62.0 60.5 61.0 60.8 59.6
Distribution Random 62.1 60.9 62.5 62.4 60.6 59.8 59.7 60.0 59.6 60.1
Appendix C Ablation of Activation Patterns for Sparse Vector $Z_{k}^{u}$
Although we have conducted experiments on the activation rate $\beta$ of the sparse vector $Z_{k}^{u}$ (Appendix B), the activation pattern itself is still worth exploring. In designing our activation strategy, we aimed to activate different neurons according to different tasks and allow the number of activations to fluctuate slightly, mirroring the human thinking process. The simplest and most intuitive approach is to predefine an activation rate and allow it to fluctuate according to a Gaussian distribution, as illustrated in Section 3.1. However, other patterns can also adhere to this activation principle. Below are additional activation strategies explored in our experiments (assuming we need to activate $\beta\%$ of neurons and the fluctuation rate is $f\%$, where $f<\beta$):
  • Top-$2\beta$ with Random Activation: First, we select the top $2\beta\%\pm 2f\%$ of neurons, then randomly activate $\beta\%\pm f\%$ of neurons within this selected group.
  • Level Random Activation: We sort and divide all neurons into 10 levels. In each level, we randomly activate $0.1\beta\%\pm f\%$ of neurons based on task instructions.
  • Distribution Random Activation: We define a hyperparameter $m$, then activate the $\beta\%\pm f\%$ of neurons closest to $m$ (in this approach, we set $m$ to 0 or 1).
The corresponding results for reasoning segmentation are shown in Table 5. Surprisingly, we found that the specific activation strategy had little effect on the results (except the last distribution random activation strategy). Instead, the activation ratio $\beta\%$ had a more significant impact on the final outcomes, indicating that the model can consistently find the neurons it needs. Therefore, to keep the paper concise and easy to understand, we employ the simplest activation strategy in Section 3.1.